40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 30 02:57:03 2015, Jenne, Update, Modern Control, First try with PRCL ASC Wiener filtering 9x
    Reply  Tue Mar 31 22:27:43 2015, Jenne, Update, Modern Control, Preliminary PRMI angular Wiener results PRMI_31Mar2015.pdf
       Reply  Sat Apr 4 17:54:03 2015, rana, Update, Modern Control, Preliminary PRMI angular Wiener results 
       Reply  Tue Apr 21 01:17:13 2015, Jenne, Update, Modern Control, T-240 self-noise propagated through stack and pendulum T240selfnoise.pngLimits.tar.gz
Message ID: 10959     Entry time: Fri Jan 30 02:57:03 2015     Reply to this: 11186
Author: Jenne 
Type: Update 
Category: Modern Control 
Subject: First try with PRCL ASC Wiener filtering 

[Aside - How do you rotate plots in the new elog?  It's showing them correctly in the attachments list below the entry, but not in the body of the log :( ]

I tried a round of PRCL ASC Wiener filtering today, but something wasn't right.  I was able to either make the cavity motion worse, or completely throw the cavity out of lock.  Making it less noisy didn't happen.

I took only 9 minutes of data the other day, since the PRMI didn't want to stay locked while it was daytime.  So, this wasn't a whole lot to train on.  But, even so, I designed some Wiener filters.  The plots with the designs show the calculated Wiener filter ("Wiener") and the result from vectfit ("Fit"). Below the bode plot is the coherence between that witness (seismometer direction) and the degree of freedom (QPD channel).  The fits were weighted by the choherence, so you can see that in the areas where the coherence was good, the fit was good.  Elsewhere, it's not so great.

 

Using these filters, and assuming a Cheby1 2nd order lowpass filter at 30Hz, I predicted the following residuals:

After discovering that these filters didn't work, I went rogue and also put in a high pass filter at 0.1 Hz, but that didn't really change anything.

Here is a plot of what happened in Yaw.  The Wiener filters' gains were all 0.3 here, which made the cavity motion larger, but not so large that it lost lock.  The filters ought to have gains of 1 - the Wiener calculation should figure out the gains appropriately, if I've given it enough information.  Here, as in the prediction plots above, red is the reference with the Wiener off, and black is with the Wiener filters on.  Black is supposed to be below red, if the filters are working.  Blue is the estimate of the angular motion that is being fed forward to the PRM, and you can see that at least the general shape is correct.  I do need to figure out what the resonance in the blue trace is from - it's at the same frequency as a peak in the T-240's spectrum (that I didn't save).  I suspect the cable might be touching the spaghetti pot on the inside, and making a mechanical short to pot vibrations.

Anyhow, more work to be done.  I left the PRMI locked for a while this afternoon, starting at 5:15ish, so I'll see tomorrow how long of a lock stretch I was able to capture for training. 

 

Attachment 1: WienerFilter_Witness1.png  438 kB  | Hide | Hide all
WienerFilter_Witness1.png
Attachment 3: WienerFilter_Witness3.png  443 kB  | Hide | Hide all
WienerFilter_Witness3.png
Attachment 9: 29Jan2015_Yaw_didnotwork.pdf  47 kB  | Hide | Hide all
29Jan2015_Yaw_didnotwork.pdf
ELOG V3.1.3-